A man in a hot-air balloon drops an apple at height of 150m. If the said balloon is

rising at 15m/s, find the highest point reached by the apple.


1
Expert's answer
2021-03-12T09:18:21-0500

Let's write the kinematic equation for the vertical displacement of the apple:


"y=y_0+v_0t-\dfrac{1}{2}gt^2,""y=150+15t-4.9t^2."

The highest point reached by the apple will be when "\dfrac{dy}{dt}=0":


"15-9.8t=0,""t=\dfrac{15\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=1.53\ s."

Substituting "t" into the previous equation we can find the highest point reached by the apple:


"y=150\ m+15\ \dfrac{m}{s}\cdot1.53\ s-0.5\cdot9.8\ \dfrac{m}{s^2}\cdot(1.53\ s)^2=161.5\ m."